OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

@dan815 @surjithayer @cwrw238

OpenStudy (anonymous):

simplify the equation to get sin2x-2sinx+1=0

OpenStudy (anonymous):

you know the formula for sin2x?

OpenStudy (anonymous):

I have a cheet sheet for formulas and identities but not on this computer

OpenStudy (anonymous):

sin2x =2sinx.cosx

OpenStudy (anonymous):

equation becomes 2sinx.cosx-2sinx+1=0

OpenStudy (anonymous):

i think it should be sin^2x

OpenStudy (anonymous):

Oh yes your right @surjithayer

OpenStudy (anonymous):

then its more easy i guess.

OpenStudy (anonymous):

Okay so how do I start?

OpenStudy (anonymous):

\[\sin ^{2}x-2\sin(x)+1=0\]

OpenStudy (anonymous):

replace sin x by t just for solving purpose.

OpenStudy (anonymous):

How did you get to that?

OpenStudy (anonymous):

the above equation?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

take -5 from right to left side which makes in 7sin^2x-14sinx+7=0

OpenStudy (anonymous):

Oh! I see it

OpenStudy (anonymous):

now replace sinx by t for solving. \[t ^{2}-2t+1=0\]

OpenStudy (anonymous):

now solve for t. i hope you know how to factorise.

OpenStudy (anonymous):

So i believe t would equal 1

OpenStudy (anonymous):

yea good. now, \[t=1\] \[sin(x)=1\]

OpenStudy (anonymous):

npw for what value of x sin x shouldbe equal to 1?

OpenStudy (anonymous):

Im not sure what you mean

OpenStudy (anonymous):

I may be over thinking it...

OpenStudy (anonymous):

we replaced sin x by t?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so if t=1 sin x should also be equal to 1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now for what degree sin is 1?

OpenStudy (anonymous):

let me tell you the answer. sin(90)=1 sin(x)=1 so x=90.

OpenStudy (anonymous):

or you can say (2n+1)*90

OpenStudy (anonymous):

OH, i did over think it ! Great, I got it. Thanks!!

OpenStudy (anonymous):

anytym.